ULTRA SENSITIVE CAVITY RING DOWN SPECTROSCOPY OF MAJOR ATMOSPHERIC SPECIES BE-TWEEN 1.20 AND 1.71 μm

\underline{A. CAMPARGUE}, S. KASSI, and D. MONDELAIN, Université Grenoble 1/CNRS, UMR5588 LIPhy, Grenoble, F-38041, France.

During recent years, we have developed a fibered DFB laser CW-CRDS spectrometer providing routine noise equivalent absorption of α_{min} ≈ 5 × 10^{−11}cm^{−1}, over the 5850-8350 cm^{−1} range. A detection limit of α_{min} ≈ 5 × 10^{−13}cm^{−1} has been recently achieved by averaging spectra over a small spectral interval. The performances of this set up have allowed extending significantly the knowledge of the absorption spectra of molecules of major importance: methane, oxygen, water, ozone, carbon dioxide, hydrogen, nitrogen. The most striking results will be presented.